OpenStudy (anonymous):

(sin 60 degree)^2(sec 240 degree)+(tan 45 degree)(cos 225 degree)^2 = ?

OpenStudy (anonymous):

\[\sin60 = \frac{\sqrt{3}}{2}\] \[\sec(240) = \sec(180 + 60) = -\sec60 = -2\] \[\tan45 = 1\] \[\cos(225) = \cos(180 + 45) = -\cos45 = \frac{-1}{\sqrt{2}}\] Put these values in the above expression and solve to get your answer..

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

\[(\sin60)^2.(\sec240) + (\tan45).(\cos225)^2\] I have provided you the values just plug in and try to solve further..

OpenStudy (anonymous):

you basically just wrote it again. i dont know how to solve it, so i need help solving it.

OpenStudy (anonymous):

What is the values of sin60?? Just reply..

OpenStudy (anonymous):

square root 3/2

OpenStudy (anonymous):

Yes.. Plug in the question this value.. Now tell me what is value of sec240??

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Like wise I have given all the values.. So, what you have to do is to plug in these values in the question.. What is problem in that?

OpenStudy (anonymous):

i dont know how to add fractions.

OpenStudy (anonymous):

Just plug in the values and then show me.. I will help in solving further..

OpenStudy (anonymous):

especially when there is square root

OpenStudy (anonymous):

I will tell you the whole thing but show me your work after plugging in the values..

OpenStudy (anonymous):

square root 3/2 * -2

OpenStudy (anonymous):

Further??

OpenStudy (anonymous):

See there is square of sin60 term.. so what will be the square of root(3)/2??

OpenStudy (anonymous):

square root 6/4 ?

OpenStudy (anonymous):

Do you know how to square??

OpenStudy (anonymous):

plain numbers yes, square roots, no!

OpenStudy (anonymous):

Use the following formula and tell me then the square of root(3): \[\sqrt{a} \times \sqrt{a} = (\sqrt{a})^2 = a\] Tell me now the square of root(3)??

OpenStudy (anonymous):

SR3*SR3=(SR3)^2=9?

OpenStudy (anonymous):

Great..! What is the value of a?? See, Compare \[\sqrt{a}\] with the \[\sqrt{3}\] and tell me what is a here??

OpenStudy (anonymous):

SR3a?